Normal body temperature T in humans is within 1F OF 99F. Use absolute value to write the range as an inequality.
Answer by Theo(13342)

i velieve the equation will be abs(99 - x) <= 1

this absolute value equation translates to:

either (99 - x) <= 1 or -(99 - x) <= 1

if you multiply both sides of the second inequality by -1, then your two inequalities become:

(99 - x) <= 1 or (99 - x) >= -1

add x to both sides of each inequality; subtract 1 from both sides of the first inequality; add 1 to both sides of the second inequality.
you will get:

98 <= x or 100 >= x

98 <= x is the same as x >= 98.

100 >= x is the same as x <= 100.

you get x >= 98 and x <= 100 to get 98 <= x <= 100.

the temperature can be anywhere between 98 and 100 inclusive and be considered normal.

your solution is |99 - x| <= 1.

this can also be written as abs(99 - x) <= 1.